OpenStudy (anonymous):

Kiran drove form Tortula to Cactus, a distance of 250 miles. She increased her speed by 10 mi/h for the 360-mi trip from Cactus to Dry Junction. If the total trip took 11 hrs, what was her speed from Tortula to Cactus?

OpenStudy (kropot72):

Time taken for trip from Tortula to Cactus = 250/S hrs where S is the speed in mi/h Time taken for trip from Cactus to Dry Junction = 350/(S + 10) hrs Total time for trip = 11 hrs \[11=\frac{250}{S}+\frac{360}{S+10}\] Now you need to solve for S. Can you do that?

OpenStudy (anonymous):

yes, I can solve for s ;) Thank you for helping me!

OpenStudy (kropot72):

You're welcome :)